### v1

@@ Appendix A Original prompts in Portuguese
This appendix presents all the prompts used in this article, in Portuguese. Appendix B shows their translated version into English.
@@
Os itens identificados deve possibilitar que o usuário redija uma matéria jornalística com início, meio e fim.
@@ Appendix B Translated prompts into English
This appendix presents the translated version of the prompts shown in Appendix B.
[System] You are an assistant who analyzes articles written by the Agência Câmara, from the Chamber of Deputies. Your role is to identify the following items in the article:
@@
Disregard the names of the journalists or editors of the article. The only people of interest are those mentioned in the body of the article.
@@
       – "cargo": string, indicating the position held by the person, along with the agency, entity, or company they work for, if available
       – "opinioes": a list of strings indicating all the opinions that the person defended and that are mentioned in the text. The opinions must be listed in detail. If it is a direct quotation, the text in the list MUST BE IDENTICAL to what is contained in the article, including the quotes.
@@
The identified items should allow the user to write a journalistic article with a beginning, middle, and end.
@@
       – "cargo": string, indicating the position held by the person, along with the agency, entity, or company they work for, if available
       – "opinioes": a list of strings indicating ALL the relevant opinions on the subject that the person defended and that are mentioned in the text. The opinions must be listed in detail
– "tl_dr": a summary that will enable the user to write a journalistic article about the public hearing using the data (names and opinions) of those involved extracted from the transcript
@@
The same person may appear in both lists, but in different ways (slight variations in the name, changes in the position…)
Your goal is to create a correspondence between the two lists in the form of a dictionary, where the keys are the person’s names from list 1, and the values are the corresponding names from list 2. Only the names should be noted, not the positions. If there is no match, write ’Not identified’.
@@
The return of your analysis should always be in JSON format and will contain two properties:
@@ 1. Identify, in the text, the excerpts that can support the analyzed opinion;
The return of your analysis should always be in JSON format and will contain three properties related to the previous steps:
@@
Figure 15: Prompt 2 to detect hallucinations (translation of 8).
@@ 1. Identify, in the text excerpts, sentences that support the analyzed opinion;
The return of your analysis should always be in JSON format and will contain three properties related to the previous steps:

### v2

@@ Appendix A Original prompts in Portuguese
This appendix presents all the prompts used in this article, in Portuguese. Appendix B contains their English translations.
@@
Os itens identificados devem possibilitar que o usuário redija uma matéria jornalística com início, meio e fim.
@@ Appendix B Translated prompts into English
This appendix presents the English translations of the prompts shown in Appendix A.
[System] You are an assistant who analyzes news articles written by the Agência Câmara, from the Chamber of Deputies. Your role is to identify the following items in the article:
@@
Ignore the names of the journalists or editors of the article. Only those mentioned in the body of the article should be considered.
@@
   – "cargo": string, indicating the position held by the person, along with the agency, organization, or company where they work for, if available
   – "opinioes": a list of strings indicating all the opinions expressed by the person that are mentioned in the text. The opinions must be described in detail. In the case of a direct quotation, the text in the list MUST match the article exactly, including quotation marks.
@@
The identified items should make it possible for the user to write a well-structured journalistic article.
@@
   – "cargo": string, indicates the position the person holds, together with the agency, organization, or company where they work, if available.
   – "opinioes": a list of strings indicating ALL the relevant opinions about the subject discussed in the hearing that are are mentioned in the text. The opinions must be listed in detail
– "tl_dr": a summary that allows the user to draft a news article about the public hearing based on the participants’ names and opinions extracted from the transcript
@@
The same person may appear in both lists, but with slight differences (such as small variations in the name or position).
Your goal is to create a mapping between the two lists in the form of a dictionary, where the keys are names from list 1, and the values are the corresponding names from list 2. Write only the names, ignore the positions. If there is no match, write ’Not identified’.
@@
The output of your analysis should always be in JSON format and will contain two properties:
@@ 1. Identify, in the text, the excerpts that can support the analyzed opinion;
The output of your analysis should always be in JSON format and will contain three properties related to the previous steps:
@@
Figure 16: Prompt 2 to detect hallucinations (translation of 9).
@@ 1. Identify, in the text excerpts, sentences that support the analyzed opinion;
The output of your analysis should always be in JSON format and will contain three properties related to the previous steps:
@@
Appendix C Confusion matrices for the experiment
Table 5: Confusion matrices for automatic hallucination detection using GPT-4o mini.
Prompt 1
Valid opinions Hallucination Total
Dataset Valid opinions 3 673 61 3 734
Hallucination 181 323 504
Total 3 854 384 4 238
Prompt 2
Valid opinions Hallucination Total
Dataset Valid opinions 3 274 460 3 734
Hallucination 89 415 504
Total 3 363 875 4 238
Prompt 3
Valid opinions Hallucination Total
Dataset Valid opinions 2 405 1 329 3 734
Hallucination 39 465 504
Total 2 444 1 974 4 238
Table 6: Confusion matrices for automatic hallucination detection using GPT-4o.
Prompt 1
Valid opinions Hallucination Total
Dataset Valid opinions 3 561 173 3 734
Hallucination 165 339 504
Total 3 726 512 4 238
Prompt 2
Valid opinions Hallucination Total
Dataset Valid opinions 3 276 458 3 734
Hallucination 90 414 504
Total 3 366 872 4 238
Prompt 3
Valid opinions Hallucination Total
Dataset Valid opinions 2 922 812 3 734
Hallucination 46 458 504
Total 2 968 1 270 4 238
Table 7: Confusion matrix for automatic hallucination detection using DeepSeek-V3.
Prompt 1
Valid opinions Hallucination Total
Dataset Valid opinions 3 666 68 3 734
Hallucination 241 263 504
Total 2 968 1 270 4 238
Prompt 2
Valid opinions Hallucination Total
Dataset Valid opinions 3 585 149 3 734
Hallucination 167 337 504
Total 2 968 1 270 4 238
Prompt 3
Valid opinions Hallucination Total
Dataset Valid opinions 3 300 434 3 734
Hallucination 95 409 504
Total 2 968 1 270 4 238
Table 8: Confusion matrix for automatic hallucination detection using Sabiá-3.1.
Prompt 1
Valid opinions Hallucination Total
Dataset Valid opinions 3 677 57 3 734
Hallucination 255 249 504
Total 2 968 1 270 4 238
Prompt 2
Valid opinions Hallucination Total
Dataset Valid opinions 3 647 87 3 734
Hallucination 221 283 504
Total 2 968 1 270 4 238
Prompt 3
Valid opinions Hallucination Total
Dataset Valid opinions 3 632 102 3 734
Hallucination 208 296 504
Total 2 968 1 270 4 238